OpenStudy (anonymous):

Predict the next number in this sequence: 1, 4, 9, 16, ... A. 25 B. 32 C. 23

OpenStudy (anonymous):

4 -1 = 9 -4 = 16 - 9 = ?

OpenStudy (anonymous):

A. i will explain if you want to know how i got that.

OpenStudy (fj):

25

OpenStudy (anonymous):

yeah go ahead [:

OpenStudy (fj):

4-1=3 9-4=5 16-9=7 25-16=9 they are rise 2 number

OpenStudy (anonymous):

oh i get ittt!!

OpenStudy (anonymous):

2 is added to each nuber in the sequence

OpenStudy (anonymous):

cool i get it now, thanks guys